### v2

# Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Cameras

Abstract
The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid motion estimation. This study investigates the tremendous potential of spike cameras (a type of ultra-high-speed, high-dynamic-range camera) in PIV. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), designed specifically for highly turbulent and intricate flow fields. To aggregate motion features from the spike stream while minimizing information loss, we incorporate a Detail-Preserving Hierarchical Transform (DPHT) module. Additionally, we introduce a Graph Encoder (GE) to extract contextual features from highly complex fluid flows. Furthermore, we present a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), which provides labeled data for three challenging fluid dynamics scenarios. Our proposed method achieves superior performance compared to existing baseline methods on PSSD. The datasets and our implementation of SIV are open-sourced in the supplementary materials.
1 Introduction
The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research. This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraftKompenhans et al. (2000), optimizing vehicle designs to reduce air resistanceNakagawa et al. (2016), improving energy conversion efficiency in combustion processesChen et al. (2022), understanding the impact of blood flow dynamics on diseasesRaschi et al. (2012), and predicting the spread of water pollutionMoayeriKashani et al. (2016).
In PIV experiments or simulations, tracer particles(e.g., silver-coated particles) are introduced into the fluid to scatter optical signals. By estimating their displacements, PIV can infer the velocities of the fluid in which they are suspended.
Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetSCai et al. (2019b) and RAFT-PIVLagemann et al. (2021).
The success of optical flow networksDosovitskiy et al. (2015); Sun et al. (2018); Teed and Deng (2020) inspires the development of learning-based PIV algorithmsCai et al. (2019b, a); Lagemann et al. (2021). However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field. Although optical flow networks generally exhibit good generalization ability, the characteristics of fluid scenes cannot be ignored for further performance improvement. Turbulence is an irregular flow phenomenon occurring at high Reynolds numbers. When measuring the turbulent velocity field with PIV, this scenario is significantly different from macroscopic worlds, with the following key challenges:
(1) The fluid often moves rapidly within the field of view.
(2) Signals are only present in regions with tracer particles, and since the particle volume is small, downsampling may lead to signal loss;
(3) Unlike solids, the fluid is highly unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
These challenges often coexist in turbulence. Challenge (1) places high demands on experimental hardware, typically requiring ultra-high-speed cameras. Challenge (2) requires us to carefully address potential resolution loss in the network. Challenge (3) introduces problems to the extraction of context features. Based on these characteristics, we propose improvements in both hardware and algorithm aspects.
In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV. We argue that using spike cameras instead of CMOS sensors in fluid measurement tasks offers several advantages:
High temporal resolution (HTR). Spike cameras can achieve ultra-high temporal resolution (20000Hz), reducing the interval $\Delta t$ between the source and target time, and providing rich temporal information for accurate predictions. This allows the spike camera to cope well with Challenge (1);
High dynamic range (HDR). When particles accumulate in large numbers, local regions may become overly bright, exceeding the dynamic range of conventional cameras. Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range compared to CMOS sensors. This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated areas);
Reduced experimental complexity. Spike cameras can continuously capture images at high frame rates with exceptionally low bandwidth and eliminate the need for synchronization controllers to modulate the camera and laser.
In terms of algorithm, we introduce a specialized network, Spike Imaging Velocimetry (SIV) to estimate fluid displacement fields. We propose two modules, the Detail-Preserving Hierarchical Transform (DPHT) designed in consideration of Challenge (2), and the Graph Encoder (GE) mainly designed for Challenge (3). We also introduce a dataset called Particle Scenes with Spike and Displacement (PSSD). PSSD comprises labeled data addressing three common motion estimation scenarios: steady turbulence (Problem 1), high-speed flow (Problem 2), and high dynamic range scenarios (Problem 3). In the experimental section, we conduct comparative analyses of our method against both image-based and spike-based approaches. Our proposed SIV achieves state-of-the-art performance across all three sub-datasets. The effectiveness of the proposed modules is further demonstrated through ablation studies. The main contributions of this work are as follows:
1. Explore a feasible approach for applying spike cameras to fluid velocity measurement;
2. From the perspective of fluid dynamics, especially turbulent flows, and the characteristics of the PIV scenarios, design a suitable network structure SIV, including two new modules: DPHT and GE;
3. Propose a PIV dataset for recording continuous fluid motion with spike cameras, which includes four flow types and three challenging experimental scenarios.
2 Related Work
2.1 Spike-based Optical Flow Estimation
The spike cameraDong et al. (2021); Huang (2019) is a type of neuromorphic camera (NeurCam). Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina. Spike cameras mimic the encoding principles of the central fovea in the retina, providing continuous visual information with ultra-high temporal resolutions, high dynamic range(HDR), and significantly reduced energy consumption and latency, making them well-suited for high-speed motion scenarios. Their pixels independently accumulate photons and emit a spike when the accumulated intensity exceeds a certain threshold. This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images. For a detailed explanation of the working mechanism, refer to Dong et al. (2021). The application of spike cameras in optical flow estimation provides significant insights for this study. Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of Particle Image Velocimetry (PIV). The techniques employed in PIV for image registration and displacement estimation are essentially a variant of optical flow estimation. Spike-based optical flow estimation is still a relatively new field. Hu et al. (2022) introduced a specialized neural network architecture for spike-based optical flow estimation. This structure can adaptively eliminate motion blur in spike streams based on prior motion information. A framework based on Teed and Deng (2020) is proposed by Zhao et al. (2022) to handle continuous spike streams and is then extended by Zhao et al. (2024) to suppress the adverse effects of randomness in spike streams. Chen et al. (2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by Xia et al. (2024).
2.2 Image-based PIV Algorithms
Many traditional image-based PIV algorithmsLourenco and Krothapalli (1995); Nobach and Honkanen (2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation areas (IA). Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent. However, traditional methods depend on finely tuned parameters and complex optimization techniques. With advancements in optical flow networks, various architecturesDosovitskiy et al. (2015); Sun et al. (2018); Teed and Deng (2020) are developed to predict dense displacement fields of macroscopic objects with high accuracy. Early PIV networks like PIV-DCNNLee et al. (2017) laid the groundwork for these developments. PIV-NetSCai et al. (2019b) is capable of generating full-resolution displacement outputs. Cai et al. (2019a) adapt LiteFlowNetHui et al. (2018) to PIV. RAFT-PIVLagemann et al. (2021), derived from RAFTTeed and Deng (2020), has shown exceptional performance on multiple datasets. Additionally, unsupervised learning frameworks for PIVZhang and Piggott (2020); Zhang et al. (2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements. Other recent PIV networks include LIMAManickathan et al. (2023), CC-FCNGao et al. (2021), En-FlowNetCJi et al. (2024a), ARAFT-FlowNetHan and Wang (2023), OFVNetSJi et al. (2024b), and PIV-PWCNetZhang et al. (2023b). However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field, and most of them have not been open-sourced.
2.3 Event-based PIV Algorithms
The event camera(such as DVSLichtsteiner et al. (2008), DAVISMoeys et al. (2017), ATISPosch et al. (2010), CeleXHuang et al. (2017)) is another type of NeurCam, which is inspired by the peripheral retinaRebecq et al. (2019). Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes. When the brightness change of a pixel exceeds a preset threshold, the camera records this change event (brightness increase as a positive event, brightness decrease as a negative event) and outputs its timestamp, position, and change value as data. However, due to the lack of absolute intensity values, event cameras struggle to reconstruct complete images, particularly in static regions and with fine textures. The application of event cameras to PIV has already been explored by Willert and Klinner (2022) and Willert (2023). Similar to the spike camera, the event camera offers higher temporal resolution, high dynamic range, low latency, and reduced experimental complexity. However, event cameras are only sensitive to objects with obvious motion and have hardly any output signal in stationary regions. This can be a weakness compared to spike cameras, considering that not all areas of the flow field may be moving at high speed.
3 Preliminary
3.1 The Working Principle of the Spike Camera
Each photosensitive element in a spike camera generates spike streams sequentially over time. A ”1” in the stream indicates that a spike has occurred (i.e., the threshold has been reached), while a ”0” indicates an ongoing accumulation state. The Integral-and-Fire (IF) mechanism of spike cameras can be mathematically represented as follows:
$A(\mathbf{x},t)=\int_{0}^{t}\alpha\cdot I(\mathbf{x},\tau)\bmod\theta,$ (1)
Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator. The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represent the time length of the spike stream. Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and brightness representation.
3.2 Data Generation
Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3). Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground truth turbulent displacement fields when $\Delta t=21$ (about 1.05 ms) and $\Delta t=11$ (about 0.55 ms). Additionally, we provide corresponding pairs of images for comparative analysis for image-based methods. We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream. The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB)Li et al. (2008). JHTDB employs Direct Numerical Simulation (DNS) to simulate turbulent flow fields. DNS directly solves the Navier-Stokes equations without any model assumptions or simplifications, enabling precise capture of all scales in turbulent flow. This method resolves the fluid motion in both space and time with high-resolution grids, accurately capturing the complex and multi-scale nature of turbulence. Additional details about the dataset can be found in our supplementary materials.
4 Approaches
4.1 Overall Architecture
Problem definition. Let $\mathbf{S}\in\mathbb{B}^{H\times W\times T}$ represent the spike stream recorded by the spike camera. $\mathbb{B}$ is the binary domain. We aim to estimate the displacement $\mathbf{D}$ of particles in the fluid between time moments $t$ and $t+\Delta t$. At the pixel level, for any given pixel $\mathbf{p}=(x,y)$ at $t$, we need to find a unique correspondence formulated by:
$\mathcal{I}(\mathbf{p},t)\leftarrow\mathcal{I}(\mathbf{p},t+\Delta t).$ (2)
Here, $\mathcal{I}$ is the mapping of the spike stream $\mathbf{S}$ in the high-dimensional feature space, representing the scene information, i.e., $\mathcal{I}=\mathrm{Rep}(\mathbf{S})$.
Network architecture. The primary structure of the SIV network is shown in Figure 2. To represent the scenes at moments $t$ and $t+\Delta t$, we extract spike stream segments (source and target) near these two moments, denoted as $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$, and input them into the network. Firstly, through the proposed DPHT module, the sub-stream $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$ are embedded into the representations $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ by our proposed DPHT module. Motion features $\mathbf{F}_{s,m}$ and $\mathbf{F}_{t,m}$ are extracted from $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$. Through the proposed Graph Encoder, the context feature $\mathbf{F}_{c}$ is extracted from $\mathbf{R}_{t}$. Subsequently, the enhanced features $\mathbf{F}_{s,m}$, $\mathbf{F}_{t,m}$ and $\mathbf{F}_{c}$ undergo an iterative optimization process in a pyramid structure, ranging from coarse (1/16 downsampled branch) to fine (1/4 downsampled branch). In each branch, similar to DIPZheng et al. (2022), the Inverse Patchmatch Block is used to construct the correlation volume, and is aggregated through a GRU. Finally, the Local Search Block estimates the residual displacement field $\Delta\mathbf{D}$ by searching for local cost based on the current displacement field estimate, and cost aggregation is performed through another GRU.
Figure 2: The overall network structure of SIV includes motion feature extraction supported by DPHT and the motion encoder, context feature extraction supported by DPHT and the Graph Encoder, along with a multi-scale iterative optimization process.
Figure 3: (a) A demonstration of the Detail-Preserving Hierarchical Transform (DPHT) module. (b) A brief illustration of Graph Encoder (GE) module.
4.2 Detail-Preserving Hierarchical Transform
Zhao et al. (2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network. Low-level features mainly focus on local details, while high-level features aggregate these low-level features to express more complex meanings. However, our ablation experiments (see Table 2) show that directly using multi-level representations does not achieve optimal performance improvement for PIV. This is because, as mentioned in the Introduction, the reduction in resolution or scale of feature maps leads to a significant loss of detail information. Furthermore, these details may not be critical for macroscopic objects, but considering that the particles in PIV are small and the fluid motion is highly intricate and complex, even the loss of detail information could result in an increase in local prediction errors. Moreover, the sparse and dark regions without particles essentially carry little to no signal, much of the information in a small region may stem from the bright particles. Based on these considerations, we designed a 3D Detail-Preserving Downsampling unit in which a bunch of 3D convolutions and 3D max poolings are incorporated (as shown in Figure 3(a)) to reduce the progressive loss of important information across the multi-layer structure while enabling the network to aggregate features at multiple scales. We refer to the entire block in Figure 3(a) as the Detail-Preserving Hierarchical Transform (DPHT).
3D Detail-Preserving downsampling. As shown in Figure 3(a), suppose the input of level $l$ is a feature vector $\mathbf{x}^{(l)}$, and the processing of $\mathbf{x}^{(l)}$ at this level is as follows:
$\mathbf{x}^{(l)}_{\mathrm{out}}=\mathscr{D}^{(l)}(\mathbf{x}^{(l)})=\mathscr{F}_{\mathrm{2C}}(\mathscr{F}_{\mathrm{3C}}(\mathbf{x}^{(l)})+\mathscr{F}_{\mathrm{3MP}}(\mathbf{x}^{(l)})),$ (3)
Here, $\mathscr{F}_{\mathrm{3C}}$ represents a 5×3×3 3D convolution, $\mathscr{F}_{\mathrm{3MP}}$ is a combination of a 3D max pooling operation and a 5×1×1 3D convolution, and $\mathscr{F}_{\mathrm{2C}}$ refers to a sequential stack of 2D convolutions. The output of the previous level serves as the input to the next level, i.e., $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}_{\mathrm{out}}=\mathscr{D}^{(l)}(\mathbf{x}^{(l)})$.
Global temporal aggregation. In this step, the output feature vectors from each level are further refined by several convolutions (for detailed information, refer to the supplementary materials), and then concatenated together to form the final representation of the DPHT:
$\mathbf{R}=\mathscr{A}_{m}[\mathbf{x}^{(l=1)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=2)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=3)}_{\mathrm{out}}],$ (4)
$\oplus$ represents concatenation, and $\mathscr{A}_{m}$ denotes convolution. As shown in Figure 3, we use DPHT as a pre-representation method before motion feature extraction and context information extraction, obtaining $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ for source sub-stream and target sub-tream, respectively.
4.3 Graph Encoder
Turbulent motion is highly complex, and extracting context information from turbulence is challenging. Therefore, Yu et al. (2021) argued that context extractors are unnecessary for PIV and removed this structure from their network. However, we hold a different view: despite the spatially chaotic nature of turbulence, there still exist multi-scale vortex structures (referred to as quasi-order structures), making the extraction of context information meaningful. To achieve this, we introduce Graph Convolutional Unit (GCU) in the context extractor, referring to the entire context extractor as the Graph Encoder. Graph convolutions weight the relationships between nodes using an adjacency matrix, performing local aggregation. Each node not only focuses on its own features but also passes information from neighboring nodes (as defined by the adjacency matrix) to itself. We believe that this approach is particularly suited for capturing spatial or topological structures in the PIV data, especially when the quasi-order structures of the flow is highly complex topologically. This situation arises particularly in cases with extremely high flow speeds (meaning high Reynolds numbers). When the Reynolds number is sufficiently high, turbulence experiences a cascade process from large-scale vortices to small-scale vortices (Kolmogorov scale) Kolmogorov (1991). This cascade process is the transfer of energy from large vortices to small vortices in turbulence, with each vortex disturbing the surrounding flow field, further complicating the flow.
Graph projection. The first step is to map the preliminary context information $\mathbf{R}_{c}=\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the graph space. The projection operation decouples the position information from the original grid features, making the resulting low-dimensional node feature representations more compact and expressive. The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}$, contain appearance features related to the shape and regional information of the scene context. Given $\mathbf{R}_{c}$, we can use the projection function $\mathscr{P}_{R\to V}(\cdot)$ to assign features with similar representations to the same node, resulting $\mathbf{V}_{c}\in\mathbb{R}^{C\times K}$, where $C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes. Since $\mathscr{P}_{R\to V}(\cdot)$ we define is a linear combination of feature vectors within the grid space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{c})$, the resulting nodes can theoretically aggregate both long-range and short-range information across the entire feature map, which could be beneficial for turbulent scenes with multi-scale quasi-order structures. This is formulated by,
$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{c,j}).$ (5)
Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{c})\in\mathbb{R}^{(H\times W)\times K}$.
Graph convolution. After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the adjacency matrix $\mathbf{A}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors. We utilize the graph convolution $\mathscr{G}$ from Selvan et al. (2020) to further propagate information across the graph. Specifically, for a single graph convolution with its parameter $W_{g}$, the operation is defined as
$\mathbf{V_{c}^{{}^{\prime}}}=\mathscr{G}(\mathbf{V}_{c},\mathbf{A})=\sigma(\mathbf{A}\mathbf{V}_{c}^{T}W_{g}),$ (6)
where $\sigma(\cdot)$ is a non-linear activation function. Two serial graph convolutional networks (GCNs) are used to capture the relationships between nodes. The node features updated by the GCNs are given by
$\mathbf{V_{c}^{\mathrm{out}}}=\mathscr{G}_{2}[\mathscr{G}_{1}(\mathbf{V}_{c},\mathbf{A}),\mathbf{A}],$ (7)
Graph re-projection. Our graph re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}=[z_{ij}]$ as inputs, generating a 2D feature map $\widetilde{\mathbf{R}}_{c}$. Since inverting the projection weight matrix is not feasible, we follow the approach in Li and Gupta (2018); Luo et al. (2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$. This is given by,
$\widetilde{\mathbf{R}}_{c}=\mathscr{P}_{V\to R}(\mathbf{V_{c}^{\mathrm{out}}},\mathbf{Z})=\mathbf{Z}\mathbf{V_{c}^{\mathrm{out}}}^{T}.$ (8)
Thus, $\mathscr{P}_{V\to R}$ thus linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters. Note that even if two pixels are assigned to the same node, they will have different features after re-projection. Therefore, the GCU may preserve the spatial details of the signal. Finally, these projection results can be integrated into an existing residual extractor network $\mathscr{R}$ introduced by Teed and Deng (2020). Final context feature $\mathbf{F}_{c}$ given by,
$\mathbf{F}_{c}=\mathscr{R}(\mathbf{R}_{c}+\alpha\widetilde{\mathbf{R}}_{c}),$ (9)
where $\alpha$ denotes a learnable parameter that is initialized as 0 and gradually performs a weighted sum.
5 Experimental Results
In this section, we evaluate our method on the PSSD dataset and compare its performance with other traditional and learning-based methods. Additionally, we conduct a series of ablation experiments to assess the effectiveness of the proposed modules.
Figure 4: Visualization examples of different algorithms in Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3) are shown. For each problem, the red box in the top optical flow map and the yellow box in the bottom EPE error map indicate the same region of the flow field.
5.1 Experimental Details
SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU. We employ the Adam optimizer with a warm-up strategy during training. The model parameters for DIP are consistent with those specified in the original paperZheng et al. (2022). For each problem, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs. The number of iterative refinements is set to 10, and the batch size is 4.
Following Cai et al. (2019b) and Lagemann et al. (2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms. Additionally, given the absence of a dedicated outlier metric in PIV, we introduced Root Mean Squared Error (RMSE) as a supplementary metric due to its sensitivity to outliers.
Since SIV outputs a prediction sequence of length $N$ at each iteration, the loss function is defined as:
$loss=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1},w_{i}=\gamma^{N-i-1}.$ (10)
We set $\gamma=0.8$ during training.
Methods Channel Isotropic Mhd Mixing Average
$\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$
Problem 1 SIV (ours) 0.242 / 0.284 0.157 / 0.182 0.918 / 1.288 0.574 / 0.765 0.638 / 0.934 0.407 / 0.563 0.700 / 1.076 0.433 / 0.616 0.625 / 0.895 0.393 / 0.532
RAFT-PIV - Image 0.377 / 0.437 0.215 / 0.249 2.181 / 2.769 1.252 / 1.574 1.336 / 1.701 0.842 / 1.072 1.875 / 2.486 1.064 / 1.394 1.442 / 1.848 0.843 / 1.072
RAFT-PIV - Spike 0.347 / 0.401 0.208 / 0.240 1.223 / 1.570 0.755 / 0.942 0.857 / 1.136 0.557 / 0.731 0.958 / 1.288 0.581 / 0.755 0.846 / 1.099 0.525 / 0.667
PIV-NetS - Image 0.545 / 0.904 0.301 / 0.478 2.560 / 3.136 1.497 / 1.819 1.574 / 2.121 0.985 / 1.284 2.247 / 2.845 1.320 / 1.651 1.731 / 2.253 1.026 / 1.308
PIV-NetS - Spike 0.553 / 0.821 0.289 / 0.424 1.507 / 1.991 0.874 / 1.124 1.252 / 1.782 0.734 / 1.019 1.382 / 1.903 0.782 / 1.040 1.173 / 1.624 0.670 / 0.902
SCFlow 0.354 / 0.468 0.222 / 0.284 2.232 / 2.706 1.305 / 1.628 1.374 / 1.703 0.901 / 1.110 2.132 / 2.684 1.169/ 1.435 1.522 / 1.890 0.899 / 1.114
HiST-SFlow 0.369 / 0.424 0.224 / 0.256 1.271 / 1.614 0.796 / 0.980 0.964 / 1.250 0.611 / 0.784 1.029 / 1.354 0.637 / 0.809 0.908 / 1.160 0.567 / 0.708
coarse-to-fine HS 6.790 / 8.448 0.509 / 0.655 7.415 / 9.354 1.998 / 3.079 7.683 / 10.783 2.558 / 4.463 9.388 / 7.274 1.954 / 3.146 7.819 / 9.751 1.755 / 2.836
Problem 2 SIV (ours) 0.366 / 0.443 0.256 / 0.300 1.721 / 2.602 1.076 / 1.521 1.098 / 1.671 0.717 / 1.026 1.482 / 2.493 0.907 / 1.412 1.167 /1.802 0.739 / 1.065
RAFT-PIV - Image 0.665 / 0.785 0.368 / 0.434 4.730 / 5.952 2.635 / 3.297 2.320 / 2.906 1.366 / 1.707 4.305 / 5.572 2.368 / 3.065 3.005 / 3.804 1.684 / 2.126
RAFT-PIV - Spike 0.688 / 0.792 0.412 / 0.477 2.410 / 3.169 1.483 / 1.882 1.468 / 1.936 0.959 / 1.236 2.071 / 2.864 1.241 / 1.649 1.659 / 2.190 1.024 / 1.311
PIV-NetS - Image 1.163 / 1.896 0.683 / 1.029 4.866 / 5.944 2.830 / 3.409 2.673 / 3.719 1.643 / 2.207 4.347 / 5.497 2.500 / 3.095 3.262 / 4.264 1.914 / 2.435
PIV-NetS - Spike 0.887 / 1.552 0.540 / 0.836 3.172 / 4.155 1.818 / 2.315 2.392 / 3.400 1.410 / 1.963 3.067 / 4.154 1.744 / 2.316 2.379 / 3.315 1.378 / 1.857
SCFlow 0.694 / 0.716 0.403 / 0.486 4.689 / 5.975 2.613 / 3.156 2.387 / 3.055 1.391 / 1.697 4.326 / 5.262 2.380 / 3.123 3.024 / 3.752 1.696 / 2.116
HiST-SFlow 0.556 / 0.641 0.364 / 0.419 2.212 / 2.947 1.365 / 1.756 1.404 / 1.879 0.998 / 1.308 1.911 / 2.693 1.179 / 1.595 1.521 / 2.040 0.976 / 1.269
coarse-to-fine HS 20.691 / 20.993 8.244 / 9.628 17.164 / 19.627 8.051 / 9.944 17.469 / 21.594 8.170 / 11.252 17.131 / 19.572 7.936 / 9.970 18.114 / 20.447 8.100 / 10.199
Problem 3 SIV (ours) 0.241 / 0.286 0.155 / 0.181 0.903 / 1.282 0.564 / 0.761 0.621 / 0.916 0.393 / 0.542 0.698 / 1.106 0.437 / 0.634 0.616 / 0.897 0.389 / 0.530
RAFT-PIV - Image 0.447 / 0.523 0.244 / 0.286 2.694 / 3.334 1.495 / 1.847 1.447 / 1.808 0.888 / 1.104 2.311 / 2.948 1.271 / 1.613 1.725 / 2.153 0.974 / 1.213
RAFT-PIV - Spike 0.320 / 0.373 0.200 / 0.232 1.203 / 1.558 0.740 / 0.939 0.798 / 1.061 0.530 / 0.698 0.946 / 1.297 0.576 / 0.770 0.817 / 1.072 0.512 / 0.660
PIV-NetS - Image 0.667 / 0.991 0.343 / 0.513 2.786 / 3.357 1.602 / 1.924 1.641 / 2.164 0.956 / 1.258 2.393 / 2.988 1.378 / 1.709 1.872 / 2.375 1.070 / 1.351
PIV-NetS - Spike 0.476 / 0.789 0.247 / 0.406 1.420 / 1.914 0.856 / 1.134 1.157 / 1.695 0.716 / 1.021 1.342 / 1.891 0.774 / 1.063 1.099 / 1.572 0.648 / 0.906
SCFlow 0.462 / 0.743 0.269 / 0.433 1.850 / 2.206 0.837 / 1.008 1.300 / 1.915 0.774 / 1.142 1.972 / 2.618 0.769 / 1.040 1.396 / 1.871 0.662 / 0.905
HiST-SFlow 0.329 / 0.381 0.180 / 0.210 1.179 / 1.523 0.640 / 0.828 0.832 / 1.093 0.452 / 0.616 0.960 / 1.302 0.476 / 0.663 0.825 / 1.075 0.437 / 0.579
coarse-to-fine HS 7.908 / 8.946 1.044 / 1.648 7.924 / 9.634 2.706 / 3.816 8.345 / 11.002 3.130 / 4.828 7.854 / 9.668 2.716 / 3.903 8.008 / 9.813 2.399 / 3.549
[…]
5.2 Comparative Experiments
To evaluate the effectiveness of our proposed method, we conduct extensive experiments on the PSSD dataset. The experiments are performed under identical settings to ensure fairness in the results.
Index Settings Problem 1 Problem 2 Problem 3
DPHT GE $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$
(A) 0.625 0.393 1.167 0.739 0.616 0.389
(B) ✓* 0.658 0.395 1.249 0.764 0.627 0.389
(C) 0.757 0.403 1.278 0.777 0.630 0.391
(D) ✓* 0.645 0.396 1.242 0.752 0.619 0.386
(E) 0.701 0.398 1.249 0.793 0.659 0.401
(F) 0.805 0.431 1.349 0.887 0.725 0.413
Table 2: Ablation study of our proposed modules, showing results for End-Point Error (EPE). Blocks colored greener indicate better performance, characterized by lower EPE values. ✓ indicates the presence of the module, ✗ indicates the absence, and ✓* indicates the removal of the submodules under this module, for DPHT, it refers to 3D-DPD, and for GE, it refers to GCU (including the projection and re-projection). Setting (A) represents our final model.
We compare our method with multiple open-source baselines, including a traditional PIV algorithmCai et al. (2018), PIV networks (PIV-NetSCai et al. (2019b), RAFT-PIVLagemann et al. (2021)), as well as their spike versions, and spike-based optical flow networks (SCFlowHu et al. (2022), HiST-SFlowZhao et al. (2024)).
Results on Problem 1. In the steady flow scenario, the fluid moves at a moderate speed. In both the time intervals of 1 and 2, SIV achieves the smallest EPE and RMSE, providing the optimal estimation. Compared to the suboptimal methods, SIV reduces the EPE by 26.1% when $\Delta t=21$. The traditional coarse-to-fine HSCai et al. (2018) method can only yield relatively reasonable predictions when the time interval and displacement are small, which reflects, to some extent, the challenge of the dataset. Figure 4 shows the displacement field predictions (optical flow maps) and error maps for different methods in PIV. The red boxes in the optical flow map correspond to the yellow boxes in the error map. The marked region in Problem 1 is a smooth boundary between the two components. By comparing with the ground truth optical flow map, it is clear that the optical flow map shapes predicted by SIV and the spike version of RAFT-PIVLagemann et al. (2021) are mostly accurate, while other methods show varying degrees of distortion in this region. In the error map, it is more intuitive to see that the red areas (high error regions) in SIV are much fewer.
Results on Problem 2. Problem 2 represents high-speed scenarios with greater particle displacements. SIV maintains its lead in all test scenarios, with a 23.3% reduction in EPE at $\Delta t=21$ compared to the suboptimal method. The traditional method struggles to make effective predictions in this scenario. In the visualization example of Problem 2 shown in Figure 4, the marked area includes a fine flow field shaped like a ”human hand” and a ”laminar flow region” with very high flow speed. The predicted flow field shape of SIV is very close to the ground truth, and the error in the high-speed ”laminar flow region” is significantly smaller than that of other methods.
Results on Problem 3. Problem 3 reflects scenarios with uneven lighting, where areas may exist that cause overexposure in traditional cameras (this issue does not occur in spike cameras due to their principle). SIV still achieves the best performance across all sub-datasets. In the visualization example of Problem 3 (see Figure 4), only SIV predicts a flow field in the marked region that is nearly consistent with the ground truth, and has the fewest red areas (high error regions) in the error map. As we anticipated, in Problem 3, spike-based methods outperform image-based methods, reflecting the high dynamic range advantage of spike cameras under uneven lighting conditions.
5.3 Ablation Study
We conduct a series of ablation experiments to evaluate the contribution of each component to overall performance, as detailed in Table 2. Each column in the table is color-coded with a gradient from yellow to green, where the greener color represents better performance. These results underscore the effectiveness of the DPHT and Graph Encoder (GE) and their submodules. In Problem 1, both DPHT and GE exhibit notable improvements. In high-speed motion scenarios (Problem 2), particles in the fluid exhibit greater displacements between the source time and target time, increasing the difficulty of accurate flow prediction. Under these circumstances, the enhancements provided by the 3D-DPD submodule in DPHT and the GCU submodule in GE (7.0% and 6.4%, respectively) are the most significant across the three Problems. In Problem 3, the distinctive brightness distribution of the scene might aid feature extraction, and the improvements from DPHT and GE are comparatively less pronounced but still remain observable.
6 Conclusions
We start from the perspective of fluid (primarily turbulent) scenarios and conduct a detailed analysis of their prominent characteristics, which may pose challenges to PIV hardware and deep learning algorithms. Based on these analyses, we explore a novel approach to estimate fluid motion using spike cameras. Additionally, we develop a tailored network architecture called SIV, designed to solve PIV problems while adapting to scenario-specific features, which integrates two novel modules: DPHT and GE. DPHT is mainly used to aggregate motion features from the spike stream while retaining scene details. GE is primarily used to extract contextual features from highly unstructured turbulent scenes. We also propose a more challenging and realistic PIV dataset, PSSD, which features multiple scenarios and captures the temporal dynamics of flow field variations. Through comparative experiments, we validate that our method outperforms existing open-source PIV algorithms and spike-based optical flow algorithms on the PSSD dataset. In ablation studies, we gradually remove the introduced modules and their submodules to prove their effectiveness and analyze their performance in different scenarios. In summary, this study explores the great potential of spike cameras in PIV and advocates for the use of customized network solutions for fluid motion estimation tasks.